2 3/7 is the same as:

a. 2×7+3

b. 17÷7

c. 2.37

4. 237%


1
Expert's answer
2021-05-10T17:19:55-0400

"2\dfrac{3}{7} = \dfrac{(7\times 2)+3}{7} = \dfrac{17}{7} = 2.428"